OpenStudy (anonymous):

if cosx = 2/3 and x is in quadrant 4, then sin x/2=?

OpenStudy (anonymous):

answer choices: A) 1/3 B) Rad (1/6) C) -1/3 D)-Rad(1/6)

OpenStudy (anonymous):

first you need \(\sin(x)\)

OpenStudy (anonymous):

do you know how to find it? in other words: if you know \(\cos(x)=\frac{2}{3}\) can you find \(\sin(x)\) ?

OpenStudy (anonymous):

/\[\sqrt{\sqrt{5}}/3\] right?

OpenStudy (anonymous):

forget the 2nd rad on the 5

OpenStudy (anonymous):

yes, but now i just realized i steered you wrong, since \[\sin(\frac{x}{2})=\pm\sqrt{\frac{1-\cos(x)}{2}}\] so really you do not need \(\sin(x)\) for this, i was mistaken (you were right though)

OpenStudy (anonymous):

so in x would it be 4/3?

OpenStudy (anonymous):

plug in \(\frac{2}{3}\) where you see \(\cos(x)\)

OpenStudy (anonymous):

\[\sin(\frac{x}{2})=\pm\sqrt{\frac{1-\cos(x)}{2}}=\pm\sqrt{\frac{1-\frac{2}{3}}{2}}\]

OpenStudy (anonymous):

now we have some arithmetic to do , and also have to figure out if it is plus or minus since \(x\) is in quadrant 4, \(\frac{x}{2}\) is in quadrant 2, therefore \(\sin(\frac{x}{2})\) is positive

OpenStudy (anonymous):

you good with the arithmetic?

OpenStudy (anonymous):

yea thanks a bunch!

OpenStudy (anonymous):

also one more thing

OpenStudy (anonymous):

sure, columbo